If A=6i-8j , B=8i+3j and C=26i+19j determine a and b so that aA+bB+C=0

1
Expert's answer
2022-05-28T08:22:56-0400
"a(6i-8j)+b(8i+3j)+25i+19j=0"

"\left\{\begin{matrix}
 6a+8b & =-25 \\
 -8a+3b & =-19
\end{matrix}\right.""a=\frac{37}{41},\quad b=\frac{-161}{41}"